a) HairNice Production Sdn Bhd is considering the production of a new conditioning shampoo which will require the purchase of new mixing machinery. 

The machinery will cost RM375,000, is expected to have a useful life of 10 years and is expected to have a salvage value of RM50,000 at the end of 10 years. 

The machinery will also need a RM35,000 overhaul at the end of year six. 

A RM40,000 increase in working capital will be needed for this investment project. 

The working capital will be released at the end of the 10 years. 

The new shampoo is expected to generate net cash inflows of RM85,000 per year for each of the 10 years. 

HairNice Production's discount rate is 16%.


i)Compute the net present value (NPV) of the investment opportunity? 


ii) Based on your answer to (a) above, should Anita go ahead with the new conditioning shampoo?      


1
Expert's answer
2021-12-03T08:35:46-0500

a.

"NPV= [ \frac{85000}{(1+16\%)^{1}} + \frac{85000}{(1+16\%)^{2}} + \frac{85000}{(1+16\%)^{3}}+ \frac{85000}{(1+16\%)^{4}}+ \frac{85000}{(1+16\%)^{5}}+\frac{85000}{(1+16\%)^{6}} +\frac{85000}{(1+16\%)^{7}} + \frac{85000}{(1+16\%)^{8}} +\frac{85000}{(1+16\%)^{9}} +\frac{85000}{(1+16\%)^{10}} ]"


"= [ P73,275.86 + P63,168.85 + P54,455.90 + P46,944.74 + P40,469.61 + P34,887.59 + P30,075.51 + P25,927.16 + P22,351.000 + P19,268.11] - P375,000"


"= P410,824.33 \u2013 P375,000"


"= P35,824.33"


b)Based on our answer above, the net value is positive, Anita can go ahead with the new conditioning shampoo project.